OpenStudy (anonymous):

factor x^(2012) - 1 using integer coefficients (and ellipses!)

OpenStudy (mathmate):

x^(2012) - 1 =(x^1006+1)(x^1006-1) =(x^1006+1)(x^503+1)(x^503-1) =(x^2+1)(x^1004-x^1002+x^1000-...-x^2+1) (x+1)(x^502-x^501+x^500....-x+1) (x-1)(x^502+x^501+x^500+...+x^2+1)